OpenStudy (idealist10):

Verify that y1=e^(2x) and y2=e^(5x) are solutions of y"-7y'+10y=0 on (-infinity, infinity).

OpenStudy (solomonzelman):

just find the derivative (1st and 2nd) (by using the chain for exponent) and plug the solutions in

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle y_1=e^{2x}}\) ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ tell me what is: 1. \(\large\color{black}{ \displaystyle \frac{d}{dx}\left( e^{2x}\right)}\) 2. \(\large\color{black}{ \displaystyle \frac{d^2}{dx^2}\left( e^{2x}\right)}\) ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

OpenStudy (idealist10):

I got it!

OpenStudy (solomonzelman):

oh, very nice:)

OpenStudy (idealist10):

Thanks a lot!

OpenStudy (solomonzelman):

and in any case, you just need to plug those (simple) derivatives into your differential equation

OpenStudy (solomonzelman):

yw

OpenStudy (solomonzelman):

I love e^(something) btw:)